### v1

@@ Abstract
Recent advances in artificial intelligence have given rise to large language models (LLMs) that not only achieve human-like performance but also share computational principles with the brain’s language processing mechanisms. While previous research has primarily focused on aligning LLMs’ internal representations with neural activity, we introduce a novel approach that leverages explainable AI (XAI) methods to forge deeper connections between the two domains. Using attribution methods, we quantified how preceding words contribute to an LLM’s next-word predictions and employed these explanations to predict fMRI recordings from participants listening to the same narratives. Our findings demonstrate that attribution methods robustly predict brain activity across the language network, surpassing traditional internal representations in early language areas. This alignment is hierarchical: early-layer explanations correspond to the initial stages of language processing in the brain, while later layers align with more advanced stages. Moreover, the layers more influential on LLM next-word prediction—those with higher attribution scores—exhibited stronger alignment with neural activity. This work establishes a bidirectional bridge between AI and neuroscience. First, we demonstrate that attribution methods offer a powerful lens for investigating the neural mechanisms of language comprehension, revealing how meaning emerges from preceding context. Second, we propose using brain alignment as a metric to evaluate the validity of attribution methods, providing a framework for assessing their biological plausibility.
@@ 1 Introduction
Recent progress in deep learning has led to the development of autoregressive language models that excel in capturing language structures and performing human-like capabilities in various linguistic tasks [1, 2, 3, 4]. This progress poses a critical question at the intersection of artificial intelligence and neuroscience: Are the observed similarities between language models and human cognition merely superficial, or do they stem from shared underlying mechanisms?
@@
Figure 1: Approach. a, Language processing in the brain and LLMs shows striking parallels. Applying XAI methods, this study investigates how XAI can reveal insights into both systems and their interrelationship. b, To test this hypothesis, we used attribution methods as representatives of XAI. Attribution methods reveal how much each previous word in a context influences the model’s decision about the next incoming word. Using a sliding window approach, we applied these methods to continuous stories. Since the sliding windows overlap, each word in the story is processed within different windows, resulting in multiple importance scores being assigned to a single word. These scores were stored in a vector for each word, forming an attribution feature space. c, We used ridge regression to quantify how well attribution-based explanations predict fMRI brain activity recorded while participants listened to the same stories provided to the LLMs. Prediction accuracy was assessed with a brain score, defined as the Pearson correlation between the predicted and actual brain responses on held-out data. We applied this approach to attribution feature spaces generated by four distinct attribution methods (Erasure, Integrated Gradients, Gradient Norm, and Gradient $\times$ Input) across three LLMs (GPT-2, Llama 2, and Phi-2), using four stories. Each feature space was independently tested to predict the fMRI activity of participants for its corresponding story. In total, 147 participants were analyzed across all stories. Additionally, we compared these attribution-based feature spaces to widely used internal LLM representations, namely activations and attention weights, using the same stories, LLMs, and participants. In the figure: x represents the input sequence, $x_{t}$ denotes the target word (incoming word), and $x^{\prime}$ is the baseline input. f represents the LLM function, while y is the recorded brain response (BOLD signal) from a voxel, and $\hat{y}$ is the predicted response. Corr. refers to the Pearson correlation used for evaluation. The attribution methods formulas are shown in full in the Methods.
In this study, we applied four types of attribution methods to three LLMs and constructed feature representations to model fMRI activity recorded during naturalistic story listening. We then evaluated the predictive power of these features in capturing brain activity, comparing them to traditional internal representations such as model activations and attention weights (Fig. 1b, c). Additionally, we examined the hierarchical relationship between these explanations and the brain’s language-processing regions. To further explore the role of model depth, we employed layer conductance to generate feature representations for individual layers and assessed their alignment with specific brain regions.
@@ 2 Results
Are LLM-derived explanations effective in explaining language processing in the brain? We explore this question through an encoding modeling approach that predicts brain activity based on explanations derived from attribution methods. Specifically, we focus on explanations that clarify why an LLM makes a particular prediction, providing a transparent window into the model’s decision-making process.
@@
2.1 Large Language Models’ Explanation
In this study, we leverage feature attribution methods, a key class of XAI techniques, to gain insights into how LLMs generate predictions. These methods quantify the contribution of each token in a given context to the model’s next-word prediction, providing a fine-grained view of the model’s internal decision-making process. Feature attribution techniques have proven valuable across various applications, including improving model interpretability [22], identifying biases [23], diagnosing model errors [24], and detecting shortcut learning patterns. In this study, we employ four distinct attribution methods, each of which is described in detail below.
@@ 2.1 Large Language Models’ Explanation
We applied attribution methods to the outputs of each model using windowed portions of the stories as input. The resulting importance scores were used to construct feature spaces tailored to each attribution method. These feature spaces served as the basis for our encoding modeling approach, enabling us to assess their alignment with brain activity. Further details on the attribution methods, computational process, and feature space design for the encoding model are provided in the Methods.
2.2 Language Models’ Explanation Predicts Brain Activity
@@
2.3 Large Language Models’ Explanation Outperforms Internal Representations in Language-Related Regions
@@
To test this hypothesis, we independently input the encoder with activations, attention, and attribution from each LLM to predict fMRI brain activity. To summarize the overall performance of each feature space, we aggregated brain scores across all LLMs and stories and identified voxels where brain scores were significantly greater than zero.
@@
2.4 Language Models’ Explanation Reveals the Hierarchy of Language Processing in the Brain
@@
We assessed hierarchical alignment by identifying each brain voxel’s preferred layer—the model layer whose conductance yielded the highest brain score (Fig. 4a). To examine how this preference varies across the brain, we analyzed the distribution of preferred layers within language-related regions. This analysis uncovered a striking hierarchical pattern: voxels in lower-level language areas, such as the auditory cortex and STG, were best predicted by early layers of GPT-2. As regions became increasingly distant from the auditory cortex and engaged in higher-order language processing, we observed stronger alignment with deeper model layers, particularly in the IFG and angular gyrus (Fig. 4b). Notably, while previous findings have identified later model layers as preferred for most brain voxels [8, 9, 47, 5], our explanation-based approach demonstrates that both early and late layers contribute significantly to brain alignment, underscoring the functional importance of initial model layers.
@@
To rule out the possibility that layer importance patterns were merely driven by word frequency or specific linguistic categories, we clustered words based on their part-of-speech (POS) tags and examined layer importance within each category. The results remained consistent: across most POS categories, the first and last layers exhibited the highest importance (Fig. 4d), indicating that their contributions extend beyond specific word types and reflect a more general role in language processing. Interestingly, early GPT-2 layers showed a stronger influence on content words, whereas function words were predominantly influenced by later layers. Among the middle layers, layer 5 emerged as particularly important, especially for function words. This pattern aligns with prior findings on layer-specific processing in LLMs, suggesting that early layers play a crucial role in structuring linguistic content [44], while later layers integrate broader contextual dependencies [48].
@@ 3 Discussion
In this study, we introduce explanations from LLMs as a novel tool to bridge the gap between language processing in the brain and artificial intelligence. Our analyses across multiple attribution methods and LLMs reveal that gradient-based attribution methods provide a robust foundation for modeling brain activity during natural language comprehension. Compared to traditional internal representations like activations and attention, attribution-based explanations demonstrate stronger alignment with early-stage language processing regions in the brain. Furthermore, we observe a hierarchical correspondence, where attributions from different model layers predict brain regions involved in parallel stages of language processing. These findings not only advance our understanding of the neural mechanisms underlying language but also establish a new framework for evaluating the plausibility of explanation methods, offering broad contributions to both neuroscience and artificial intelligence.
Attribution vs internal feature representation
To interpret these results from a neuroscientific perspective, it is crucial to consider the fundamental differences between attribution-based explanations and previously used internal representations.
@@
Third, attribution-based explanations inherently encode information about the next actual word in the sequence, often referred to as the target word. In contrast, internal representations contain a mixture of predictive and non-predictive information, making them more generalizable but less specific to the model’s next-word prediction task. This distinction is crucial because it positions attribution methods as a more direct tool for studying predictive mechanisms in the brain, addressing key concerns raised about previous studies that relied solely on internal representations.
Predictive coding and LLMs
A well-established claim in prior literature is that language models that achieve better next-word prediction performance tend to show stronger alignment with brain activity, which has been interpreted as evidence for predictive coding in the brain. However, this interpretation has been challenged by a recent critique from Antonello et al. [17], which raises two primary concerns. One issue is that internal representations are too general to serve as definitive evidence of predictive processing, as they encode a broad range of linguistic features beyond just prediction. Additionally, when comparing the brain alignment of layers to their next-word information (measured through perplexity), a mismatch emerges: brain alignment peaks in the middle layers, while next-word information is strongest in the final layers.
Our findings provide a compelling response to these concerns. Unlike internal representations, attribution-based explanations are inherently task-specific, as they quantify the contribution of each input word to predicting the target word. Moreover, Antonello et al. [17] used perplexity as a measure of predictive information, assuming that layers with greater predictive information should also exhibit stronger brain alignment. However, this assumption overlooks the fact that a layer’s performance in brain prediction extends beyond the predictive information within its representations. A more precise approach is required—one that captures the causal influence of each layer on next-word prediction.
To address this, our study introduces layer conductance as a more robust alternative to perplexity, as it directly measures how much each layer influences the model’s prediction. Applying this method, we find that the layers most critical for next-word prediction—identified through layer conductance—are also the ones that best align with brain activity across a higher proportion of voxels. By adopting the same reasoning as Antonello et al. [17] but using a more reliable metric, we demonstrate that the layers most causally important for prediction are also the most brain-like. This directly counters the second concern raised by the critique and reinforces the idea that predictive processing in the brain follows a hierarchical pattern similar to that observed in LLMs. Together, our findings establish attribution-based explanations as a more precise tool for investigating predictive coding in the brain.
Evaluating AI explainability by brain alignment
Beyond their neuroscientific implications, our findings introduce a novel evaluation framework for explainability methods in AI. Evaluating explanations remains a critical and unresolved challenge in XAI, particularly as its primary goal is to help end-users better understand AI systems and make informed decisions. Existing evaluation methods typically rely on two approaches: algorithmic-based evaluations, which assess how well explanations align with the model’s internal reasoning [50, 51, 52, 53], and human-centered evaluations, which measure how explanations affect user understanding and decision-making [54, 55, 56, 57]. While algorithmic approaches benefit from established metrics like faithfulness, human-centered evaluations often suffer from biases such as confirmation bias and label leakage [58], as well as practical limitations like the need for controlled experiments and participant training. Our findings suggest that brain alignment can serve as an alternative, objective evaluation criterion for explanations.
@@
Explanations that exhibit stronger alignment with brain activity may also be more comprehensible to users, as they reflect the cognitive processes engaged during naturalistic task performance. This brain-centered evaluation offers key advantages over traditional approaches. First, it provides ecological validity, as neural recordings can be obtained during naturalistic language processing rather than in constrained experimental settings. Second, it ensures scalability and reusability, allowing brain responses to be used for evaluating multiple XAI methods without requiring separate data collection for each explanation type. Finally, it enables objective measurement, as neural alignment provides a direct and quantifiable metric of explanation quality, free from the subjectivity of human judgments. By shifting the focus from behavioral validation to neurobiological alignment, this framework offers a scalable and unbiased approach for assessing the plausibility of XAI methods, bridging cognitive science and AI explainability research.
Future work
@@ 4.1.1 Brain Parcellation
4.2 Language Model Representations
@@ 4.2.4 Activations
4.3 Aligning Brain and Language Model Representations
@@ 4.5 Noise Ceiling
Supplementary Information
Figure S1: Brain score across attribution methods and LLMs in the right hemisphere. a, Voxel-wise brain scores for each combination of attribution method, LLM, and individual, averaged across individuals and voxels within the right hemisphere. This panel is analogous to Fig. 2b but for the right hemisphere. b, Brain scores of LLMs across right-hemisphere ROIs. Alignment is expressed as brain score normalized and presented as a percentage of the noise ceiling, estimated using intersubject correlation. This panel mirrors Fig. 2c. Markers represent the mean brain score (as % of noise ceiling) for each ROI. Error bars indicate the 95% confidence interval across individuals. Marker colors match the brain map, indicating the corresponding brain region, while error bar colors represent different LLMs.
Figure S2: Noise Ceiling Estimation. a, Noise ceiling values for each voxel, estimated as the average ISC across all individuals. b, Noise ceiling values for each ROI, calculated by averaging voxel-wise noise ceiling estimates within each ROI. Boxplots represent the distribution of averaged noise ceiling values across individuals.
Figure S3: Brain scores of feature spaces across right-hemisphere ROIs. Alignment is expressed as brain score, normalized and presented as a percentage of the noise ceiling, estimated using ISC. This figure mirrors Fig. 3b. Markers represent the mean brain score (as % of noise ceiling) for each ROI. Error bars indicate the 95% confidence interval across individuals. Marker colors match the brain map, indicating the corresponding brain region, while error bar colors represent different feature spaces.
Figure S4: Brain score obtained with attribution methods and internal representations. Scores are computed for each voxel and individual independently, and average across all individuals. Only significant predicted voxels are color coded (FDR corrected, $P<0.05$).
Figure S5: Brain scores predicted by GPT-2 layer conductance. Brain scores were computed independently for the conductance of each GPT-2 layer, for each voxel, and for each individual. The statistical test (Wilcoxon signed-rank) was applied separately for each layer, and voxels significantly predicted by at least one layer were color-coded (FDR corrected, $P<0.05$). The color bar represents the average brain score across all layers.
Figure S6: Comparing brain scores of layer conductance with attribution methods and internal representations for GPT-2 across ROIs, presented separately for the left and right hemispheres. a, Brain scores for the left hemisphere. b, Brain scores for the right hemisphere. Brain scores were normalized and presented as a percentage of the noise ceiling. Brain scores were calculated independently for each combination of feature space and individual, then averaged across individuals. For layer conductance, scores were averaged across all layers of GPT-2. Markers represent the mean brain score (as a percentage of the noise ceiling) for each ROI. Error bars indicate the 95% confidence interval across individuals. Marker colors correspond to brain regions, while error bar colors represent different feature spaces.
Figure S7: Relationship between layer-wise prediction performance and brain alignment across datasets. The percentage of brain voxels aligned with each layer (blue) was compared with the percentage of words most influenced by each layer (orange), based on conductance scores. Results are shown separately for each dataset, where each dataset corresponds to a story presented both to the language model (for next-word prediction) and to human participants (during fMRI recordings). The correlation within each dataset reflects the relationship between the importance of model layers in processing the story for next-word prediction and their predictive performance in modeling brain activity for the same story.

### v3

@@ Abstract
Large language models (LLMs) not only exhibit human-like performance but also share computational principles with the brain’s language processing mechanisms. While prior research has focused on mapping LLMs’ internal representations to neural activity, we propose a novel approach using explainable AI (XAI) to strengthen this link. Applying attribution methods, we quantify the influence of preceding words on LLMs’ next-word predictions and use these explanations to predict fMRI data from participants listening to narratives. We find that attribution methods robustly predict brain activity across the language network, revealing a hierarchical pattern: explanations from early layers align with the brain’s initial language processing stages, while later layers correspond to more advanced stages. Additionally, layers with greater influence on next-word prediction—reflected in higher attribution scores—demonstrate stronger brain alignment. These results underscore XAI’s potential for exploring the neural basis of language and suggest brain alignment for assessing the biological plausibility of explanation methods.
@@ 1 Introduction
Recent progress in deep learning has led to the development of autoregressive language models that excel in capturing language structures and performing human-like capabilities in various linguistic tasks [1, 2, 3, 4]. This progress poses a critical question at the intersection of artificial intelligence (AI) and neuroscience: Are the observed similarities between language models and human cognition merely superficial, or do they stem from shared underlying mechanisms?
@@
Figure 1: Approach. a, Language processing in the brain and LLMs shows striking parallels. Applying XAI methods, this study investigates how XAI can reveal insights into both systems and their interrelationship. b, To test this hypothesis, we used attribution methods as representatives of XAI. Attribution methods reveal how much each previous word in a context influences the model’s decision about the next incoming word. Using a sliding window approach, we applied these methods to continuous stories. Since the sliding windows overlap, each word in the story is processed within different windows, resulting in multiple importance scores being assigned to a single word. These scores were stored in a vector for each word, forming an attribution feature space. c, We used ridge regression to quantify how well attribution-based explanations predict fMRI brain activity recorded while participants listened to stories. Prediction accuracy was assessed with a brain score, defined as the Pearson correlation between the predicted and actual brain responses on held-out data. We applied this approach to attribution feature spaces generated by four distinct attribution methods (Erasure, Integrated Gradients, Gradient Norm, and Gradient $\times$ Input) across three LLMs (GPT-2, Llama 2, and Phi-2), using four stories. Each feature space was independently tested to predict the fMRI activity of participants for its corresponding story. In total, 147 participants were analyzed across all stories. Additionally, we compared these attribution-based feature spaces to widely used internal LLM representations, namely activations and attention weights, using the same stories, LLMs, and participants. In the figure: x represents the input sequence, $x_{t}$ denotes the target word (incoming word), and $x^{\prime}$ is the baseline input. f represents the LLM function, while y is the recorded brain response (BOLD signal) from a voxel, and $\hat{y}$ is the predicted response. Corr. refers to the Pearson correlation used for evaluation. The attribution methods formulas are shown in full in the Methods.
In this study, we applied four types of attribution methods to three LLMs and constructed feature representations to model fMRI activity recorded during naturalistic story listening. We then evaluated the predictive power of these features in capturing brain activity, comparing them to traditional internal representations such as model activations and attention weights (Fig. 1b, c). To investigate how attribution-based explanations vary across different model layers, we employed layer conductance to generate feature representations for individual layers and assessed their alignment with specific brain regions.
@@ 2 Results
Are LLM-derived explanations effective in explaining language processing in the brain? We explore this question through an encoding modeling approach that predicts brain activity based on explanations derived from attribution methods. Specifically, we focus on explanations that clarify why an LLM makes a particular prediction, providing a more transparent window into the model’s decision-making process.
@@
2.1 Explanations
In this study, we leverage feature attribution methods, a key class of XAI techniques, to gain insights into how LLMs generate predictions. These methods quantify the contribution of each token in a given context to the model’s next-word prediction, providing a fine-grained view of the model’s internal decision-making process. Feature attribution techniques have proven valuable across various applications, including improving model interpretability [22], identifying biases [23], and diagnosing model errors [24]. In this study, we employ four distinct attribution methods, each of which is described in detail below.
@@ 2.1 Explanations
We applied attribution methods to outputs of each model using windowed portions of the stories as input. The resulting importance scores were used to construct feature spaces tailored to each attribution method. These feature spaces formed the basis for our encoding approach, enabling us to assess their alignment with brain activity. Further details on the attribution methods, computational process, and feature space design for the encoding model are provided in the Methods.
2.2 Explanations Predict Brain Activity
@@
2.3 Explanations Outperform Internal Representations in Language-Related Regions
@@
To test this hypothesis, we independently trained the encoder with activations, attention, and attribution from each LLM to predict fMRI brain activity. To summarize the overall performance of each feature space, we aggregated brain scores across all LLMs and stories and identified voxels where brain scores were significantly greater than zero.
@@
2.4 Explanations Reveal the Hierarchy of Language Processing in the Brain
@@
We assessed hierarchical alignment by identifying each brain voxel’s preferred layer—the model layer whose conductance yielded the highest brain score (Fig. 4a). To examine how this preference varies across the brain, we analyzed the distribution of preferred layers within language-related regions. This analysis uncovered a striking hierarchical pattern: voxels in lower-level language areas, such as the auditory cortex and STG, were best predicted by early layers of GPT-2. As regions became increasingly distant from the auditory cortex and engaged in higher-order language processing, we observed stronger alignment with deeper model layers, particularly in the IFG and angular gyrus (Fig. 4b). Notably, while previous studies have shown that the best encoding performance is obtained from representations in the middle layers of language models [8, 9, 47, 5], our explanation-based approach reveals that both early and late layers contribute significantly to brain alignment.
@@
To rule out the possibility that layer importance patterns were merely driven by word frequency or specific linguistic categories, we clustered words based on their part-of-speech (POS) tags and examined layer importance within each category. The results remained consistent: across most POS categories, the first and last layers exhibited the highest importance (Fig. 4d), indicating that their contributions extend beyond specific word types and reflect a more general role in language processing. Interestingly, early GPT-2 layers showed a stronger influence on content words, whereas function words were predominantly influenced by later layers. This pattern aligns with prior findings on layer-specific processing in LLMs, suggesting that early layers play a crucial role in structuring linguistic content [44], while later layers integrate broader contextual dependencies [48].
@@ 3 Discussion
In this study, we introduce explanations from LLMs as a novel tool to bridge the gap between language processing in the brain and computational models. Our analyses across multiple attribution methods and LLMs reveal that gradient-based attribution methods provide a robust foundation for modeling brain activity during natural language comprehension. Compared to traditional internal representations like activations and attention, attribution-based explanations demonstrate stronger alignment with early-stage language processing regions in the brain. Furthermore, we observe a hierarchical correspondence, where attributions from different model layers predict brain regions involved in parallel stages of language processing.
The neuroscientific contributions of these findings emerge from the fundamental differences between attribution-based explanations and previously used internal representations.
@@
Third, attribution-based explanations inherently encode information about the next word in a sequence, as they measure how input words influence a specific target prediction. In contrast, internal representations capture both predictive and broader linguistic information—such as background knowledge and abstract reasoning [50, 51]—making them more generalizable across diverse tasks. However, this generalizability makes internal representations less tied to next-word prediction and, therefore, less reliable for studying predictive processing in the brain [17].
This limitation becomes evident when comparing brain alignment across model layers. Prior studies have shown that peak brain alignment occurs in middle-layer representations [10, 5, 9], rather than in deeper layers, which contain the highest predictive information. This suggests that strong alignment between internal representations and brain activity does not necessarily indicate predictive coding but may instead result from the greater generalizability of middle-layer representations [17]. By contrast, our layer-wise analysis using conductance reveals a strong correlation between a layer’s importance for next-word prediction and its brain alignment. This pattern arises because conductance not only captures stimulus-driven information within a layer but also quantifies its causal influence on the model’s next-word prediction. These findings establish attribution-based explanations as a more precise tool for investigating predictive coding in the brain.
Beyond their neuroscientific implications, our findings provide a new framework for evaluating explanations in XAI. A key challenge in XAI is determining how effectively these explanations improve user understanding of AI systems. This is typically assessed through behavioral tests [52, 53, 54, 55], but these tests often produce inconsistent results due to biases like confirmation bias and label leakage [56]. They also face practical limitations, such as the need for controlled experiments and participant training. Our approach suggests brain alignment as an alternative, objective evaluation criterion for explanations.
@@
Explanations that exhibit stronger alignment with brain activity may also be more comprehensible to users, as they reflect the cognitive processes engaged during naturalistic task performance. This brain-centered approach provides ecological validity through neural recordings in realistic contexts, scalability by reusing brain data across multiple XAI methods, and objectivity via a quantifiable measure of explanation quality, free from subjective human judgments. By shifting from behavioral validation to neurobiological alignment, this framework provides a scalable, unbiased method for evaluating XAI explanations, bridging cognitive science and AI explainability research.
@@ 4.1.1 Brain Parcellation
4.2 Large Language Model Representations
@@ 4.2.4 Activations
4.3 Aligning Brain and Large Language Model Representations
@@ 4.5 Noise Ceiling
Acknowledgments
We thank Richard Antonello and Alona Fyshe for their valuable discussions and constructive feedback on the manuscript.